OpenStudy (anonymous):

Can someone please tell me the steps for solving 2x^4(4x^2 + 3x + 1) ? I know that the answer is 8x^6+6x^5+2x^4

OpenStudy (anonymous):

2x^4*4x^2+2x^4+3x+2x^4 we know that x^a*x^b=x^(a+b) in this way 2*4=6 & x^4*x^2=x^8 similarly x4*x=x^5 ......,

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

And the answer is 8x^6+6x^5+2x^4 right?

OpenStudy (anonymous):

thats what I got

OpenStudy (anonymous):

@chetan552

OpenStudy (anonymous):

Theres now doubt just want to be sure! :)

OpenStudy (anonymous):

Thanks! :)